(3/(X^2-9))is equal to (6/(X+3))
Answer by arunpaul(104)

or(3/(X^2-9))= (6/(X+3))
or(3/(x+3)(x-3))=(6/(x+3))
or(3/(x-3))=6 (cancelling (x+3) from both side)
or 3= 6(x-3)
or 1=2(x-3)
or 1=2x-6
or 2x = 7
or x = 3.5 (Ans)